OpenStudy (marsxtc):

Domain and range of 2sinpietheta? and -5costheta/2? I know the domain is all real # but the range?

OpenStudy (psymon):

Do you know the range of sin and cos normally? Like if it were just sinx cosx?

OpenStudy (marsxtc):

I figured out the -5costheta/2 ( I believe) Wouldn't the range of sinx just be (-1,1)

OpenStudy (psymon):

Right, and the range is the same with cosx, [-1,1]. Now the absolute value of whatever number you have multiplying cosx or sinx also multiplies the end points of the range. For example, if I have 2sinx, the range becomes [-1*2, 1*2] = [-2,2]. Same thing with your problem (-5/2)cosx [-1*(5/2), 1(5/2)] = [-5/2, 5/2]

OpenStudy (marsxtc):

It was actually -5(costheta/2)

OpenStudy (marsxtc):

I figured out the range was -5,5 though. Thanks! Wouldn't it be the same for 2sin pie x? (-2,2)?

OpenStudy (psymon):

Ah, lol. Well, same idea though except itd just be [-1*5, 1*5]= [-5,5] And yes, exactly. Also, the number inside of the angle does nothing to its range. You can have sin(4235734657x), its still a range of [-1.1]. Oh, and those are supposed to be brackets, not parenthesis.

OpenStudy (marsxtc):

Ah yes.

OpenStudy (psymon):

So yeah, you got your answer then it seems :3

OpenStudy (marsxtc):

Is there any way you can explain brackets and parenthesis more to me?

OpenStudy (psymon):

Well, the idea is parenthesis means the number is NOT included in the domain and brackets means the number is included. -1 and 1 still are defined, so you have to use brackets.

OpenStudy (marsxtc):

So lets say if it never hit -1 or 1. It hit just -.99 and .99. Would it be (-1,1)?

OpenStudy (psymon):

Right. The -1, 1 are the boundaries of the range, its just whether or not those boundaries are defined or not. Like if we had something like 1/x. Well, the domain would be (-inf, 0) U (0, inf). In this case, I have to have parenthesis around the 0. Because the boundary of the domain is 0, but 0 doesnt count, itd make the fraction undefined. On the other hand if we had something like sqrt(x), the domain would be [0,inf) The boundary is 0 because we cannot have a negative square root, but the number 0 itself can exist, so thats why I need brackets in this case. So basically it depends on whether or not those boundary values actually can exist or not.

OpenStudy (marsxtc):

Alrighty. Thank you!

OpenStudy (psymon):

Yep. np :3